# Open Vocabulary Keyword Spotting through Transfer Learning from Speech Synthesis

### v1

@@ I Introduction
Keyword spotting (KWS) is a process of detecting specific keywords within a continuous audio stream, which is crucial for enabling voice-driven interactions on edge devices [11, 20]. Increasing demand for personalized voice assistants highlights the significance of user-defined keyword spotting, which is known as custom keyword detection or open vocabulary keyword spotting [16]. In contrast to closed vocabulary keyword spotting [17], where only predetermined keywords are recognized, custom keyword spotting, deals with the difficulty of identifying random keywords that the model might not have been exposed to during training, introducing an additional level of complexity to the task. Moreover, achieving high accuracy and minimizing latency are challenges in this task.
In the literature, there exist different methods for custom keyword spotting. One such method is the query-by-example (QbyE) [1, 6, 12] approach, which involves matching input queries with pre-enrolled examples. However, the effectiveness of the QbyE method relies heavily on the similarity between the recorded speech during enrollment and the subsequent evaluated speech recordings. Factors like diverse vocal characteristics among users and environments with background noise pose significant challenges that can impact the consistency of performance of the QbyE method. In addressing this concern, researchers have delved into text enrollment-based methods [16, 19]. One such method is the automatic speech recognition-based approach [10], which detects phonetic patterns in input streams and then compares them with enrolled keyword representation. However, the effectiveness of this method heavily depends on the accuracy of the acoustic model. In [19], an attention-based cross-modal matching approach is proposed that is trained in an end-to-end manner with monotonic matching loss and keyword classification loss. In [8], a zero-shot KWS is proposed, coupled with phoneme-level detection loss. Also, [13] introduced Dynamic Sequence Partitioning to optimally partition the audio sequence to match the length of the word-based text sequence. These recent end-to-end techniques [8, 13, 19], primarily depend on evaluating speech and text representations in a common latent space, demonstrated promising results in the custom keyword spotting task. Therefore, we considered the end-to-end approach for our investigation.
@@
The effectiveness of the end-to-end approaches relies on the accuracy of projecting audio and text representations into a shared embedding space. While these methods [8, 19] project speech and text representations into a shared latent space, they encounter difficulties in distinguishing between pairs of closely related pronunciations. This challenge can be addressed by generating text representations that have some acoustic knowledge. On the other hand, a TTS model converts text to audio, and the intermediate representations derived from this process possess an understanding of audio projections. Consequently, insights acquired from a pre-trained TTS model can serve as meaningful representations of text. In [2], a similar approach is used to transfer knowledge from a pre-trained TTS model for voice conversion task. Hence, in this paper, a novel framework is introduced to distill knowledge from a pre-trained TTS model for open-vocabulary keyword spotting task. The contributions of this study can be summarized as follows:
@@ II-A Text Encoder
It consists of a pre-trained Tacotron 2 [18] model, a recurrent sequence-to-sequence TTS system, which takes character sequence as input for the corresponding keyword. The resulting intermediate representations from the TTS model are then directed to a bidirectional gated recurrent unit (Bi-GRU) layer with a dimension of 64. The dimensions of the intermediate representations from the TTS model depend on the specific layer from which they are extracted. Further details of the intermediate representations are discussed in subsection III-B. The output from the Bi-GRU layer is then fed to a dense layer of size 128 units. The output from the text encoder is denoted as $T_{E}\in\mathbb{R}^{d\times m}$, where d and m denote the embedding dimension and length of the text (i.e. no of characters), respectively. The primary motivation for incorporating the pre-trained TTS system into the text encoder is to generate text representations that are aware of acoustic projections. This integration simplifies the task of projecting audio and text embeddings in shared latent space. Although this strategy does not explicitly involve generating audio from text, but exploits knowledge transfer from intermediate representations of TTS.
@@ II-B Audio Encoder
The input to the audio encoder is 80-dimensional mel-filterbank coefficients, extracted for every 10ms with a 25ms window length. To process this audio feature, the encoder consists of two 2-D convolution layers (Conv 2D) with a kernel size of 3. The first convolution layer consists of 32 filters, and the second convolution layer consists of 64 filters. Batch normalization is applied after each convolution layer to ensure stable training. To improve computational efficiency, the initial convolutional layer employs a stride of 2 to effectively reduce the number of processed frames by skipping consecutive frames. Following the convolution layers, two Bi-GRU layers each having a dimension of 64 are employed. Subsequently, the output from the last Bi-GRU layer is passed into a dense layer which generates a 128-dimensional audio embedding. The output from the audio encoder is denoted as $A_{E}\in\mathbb{R}^{d\times n}$, where d and n denote the embedding dimension and length of the audio (i.e. no of frames), respectively.
@@ G Q LPE LPH G Q LPE LPH
CTC [12] 31.65 18.23 14.67 35.22 66.36 89.69 92.29 69.58
Attention [1] 14.75 49.13 28.74 41.95 92.09 50.13 78.74 62.65
Triplet [16] 35.6 38.72 32.75 44.36 71.48 66.44 63.53 54.88
CMCD [19] 27.25 12.15 8.42 32.9 81.06 94.51 96.7 73.58
@@ II-C Pattern Extractor
It is based on the cross-attention mechanism [7, 21] which captures the temporal correlations between the audio and text embeddings. In this setup, the audio embedding $A_{E}$ functions as both the key and value, while the text embedding $T_{E}$ acts as the query. The output of the pattern extractor is the context vector which contains information about audio and text agreement.
@@ III-A Database
For training and evaluation, we used the Libriphrase [19] dataset, which comprises short phrases with varying word lengths (1 to 4). The dataset was generated from the Librispeech corpus [15]. The training set of Libriphrase is generated using train-clean-100 and train-clean-360 subsets, and the evaluation set is using train-others-500 subset. The evaluation set consists of 4391, 2605, 467, and 56 episodes of each word length respectively. Each episode has three positive and three negative pairs. The negative samples are further categorized into easy and hard based on Levenshtein distance [9], leading to the creation of the Libriphrase Easy (LPE) and Libriphrase Hard (LPH) datasets. Each sample in the dataset is represented by using three entities: audio, text, and a binary target value indicating 1 for a positive pair and 0 for a negative pair.
To comprehensively evaluate model performance, we extended our assessment beyond the Libriphrase dataset by incorporating the Google Speech Commands V1 dataset (G) [22] and the Qualcomm Keyword Speech dataset (Q) [4]. The Google Speech Commands V1 dataset (G) encompasses speech recordings from 1,881 speakers, focusing on 30 small keywords. On the other hand, the Qualcomm Keyword Speech dataset (Q) comprises 4,270 utterances of four keywords spoken by 50 speakers. Each speaker in this dataset contributes around 22-23 instances for each keyword.
@@ III-B Tacotron 2 Embeddings
In this study, the utilization of intermediate representation from Tacotron 2 is proposed to enhance performance, particularly in scenarios involving speech-keyword pairs with similar pronunciations. The intermediate representations are obtained using NVIDIA Tacotron 2 model [14] which is pre-trained on the LJSpeech dataset [3]. The details of different intermediate representations that are used in this study are outlined in Table II. Additional insights into the Tacotron 2 architecture can be found in [18]
@@ III-C Implementation Details
The training pipeline is structured as a binary classification task, wherein the model is tasked with classifying the similarity of input pairs {text, audio}. In audio and text encoders, weights are initialized through Xavier initialization, and Leaky ReLU is utilized as the activation function. A dropout of 0.2 is added after each layer in the encoders to prevent overfitting. The training process uses binary cross-entropy loss as the training criterion and employs the Adam optimizer [5] with default parameters for optimization. A fixed learning rate of $10^{-4}$ and a batch size of 128 is employed in the training process. The best-performing model was chosen based on its performance on the validation set. For training, we used four NVIDIA GeForce RTX 2080 Ti GPUs.

### v2

@@ I Introduction
Keyword spotting (KWS) is a process of detecting specific keywords within a continuous audio stream, which is crucial for enabling voice-driven interactions on edge devices [lopez2021deep, tang2018deep]. Increasing demand for personalized voice assistants highlights the significance of user-defined keyword spotting, which is known as custom keyword detection or open vocabulary keyword spotting [sacchi2019open, kgurugubelli2024comparative]. In contrast to closed vocabulary keyword spotting [sainath15b_interspeech], where only predetermined keywords are recognized, custom keyword spotting, deals with the difficulty of identifying random keywords that the model might not have been exposed to during training, introducing an additional level of complexity to the task.
In the literature, there exist different methods for custom keyword spotting. One such method is the query-by-example (QbyE) [huang2021query, lugosch2018donut] approach, which involves matching input queries with pre-enrolled examples. However, the effectiveness of the QbyE method relies heavily on the similarity between the recorded speech during enrollment and the subsequent evaluated speech recordings. Factors like diverse vocal characteristics among users and environments with background noise pose significant challenges that can impact the consistency of performance of the QbyE method. In addressing this concern, researchers have delved into text enrollment-based methods [sacchi2019open, shin22_interspeech]. One such method is the automatic speech recognition-based approach [liu2021neural], which detects phonetic patterns in input streams and then compares them with enrolled keyword representation. However, the effectiveness of this method heavily depends on the accuracy of the acoustic model. In [shin22_interspeech], an attention-based cross-modal matching approach is proposed that is trained in an end-to-end manner with monotonic matching loss and keyword classification loss. In [lee23d_interspeech], a zero-shot KWS is proposed, coupled with phoneme-level detection loss. Also, [nishu23_interspeech] introduced Dynamic Sequence Partitioning to optimally partition the audio sequence to match the length of the word-based text sequence. These recent end-to-end techniques [lee23d_interspeech, nishu23_interspeech, shin22_interspeech], primarily depend on evaluating speech and text representations in a common latent space, demonstrated promising results in the custom keyword spotting task. Therefore, we considered the end-to-end approach for our investigation.
@@
The effectiveness of the end-to-end approaches relies on the accuracy of projecting audio and text representations into a shared embedding space. While these methods [shin22_interspeech, lee23d_interspeech] project speech and text representations into a shared latent space, they encounter difficulties in distinguishing between pairs of closely related pronunciations. This challenge can be addressed by generating text representations that have some acoustic knowledge. On the other hand, a TTS model converts text to audio, and the intermediate representations derived from this process possess an understanding of audio projections. Consequently, insights acquired from a pre-trained TTS model can serve as meaningful representations of text. In [huang20i_interspeech], a similar approach is used to transfer knowledge from a pre-trained TTS model for voice conversion task. Hence, in this paper, a novel framework is introduced to distill knowledge from a pre-trained TTS model for open-vocabulary keyword spotting task. The contributions of this study can be summarized as follows:
@@ II-A Text Encoder
It consists of a pre-trained Tacotron 2 [shen2018natural] model, a recurrent sequence-to-sequence TTS system, which takes character sequence as input for the corresponding keyword. The resulting intermediate representations from the TTS model are then directed to a bidirectional gated recurrent unit (Bi-GRU) layer with a dimension of 64. The dimensions of the intermediate representations from the TTS model depend on the specific layer from which they are extracted. Further details of the intermediate representations are discussed in subsection III-B. The output from the Bi-GRU layer is then fed to a dense layer of size 128 units. The output from the text encoder is denoted as $T_{E}\in\mathbb{R}^{d\times m}$, where d and m denote the embedding dimension and length of the text (i.e. number of characters), respectively. The primary motivation for incorporating the pre-trained TTS system into the text encoder is to generate text representations that are aware of acoustic projections. This integration simplifies the task of projecting audio and text embeddings in shared latent space. Although this strategy does not explicitly involve generating audio from text, but exploits knowledge transfer from intermediate representations of TTS.
@@ II-B Audio Encoder
The input to the audio encoder is 80-dimensional mel-filterbank coefficients, extracted for every 10ms with a 25ms window length. To process this audio feature, the encoder consists of two 2-D convolution layers (Conv 2D) with a kernel size of 3. The first convolution layer consists of 32 filters, and the second convolution layer consists of 64 filters. Batch normalization is applied after each convolution layer to ensure stable training. To improve computational efficiency, the initial convolutional layer employs a stride of 2 to effectively reduce the number of processed frames by skipping consecutive frames. Following the convolution layers, two Bi-GRU layers each having a dimension of 64 are employed. Subsequently, the output from the last Bi-GRU layer is passed into a dense layer which generates a 128-dimensional audio embedding. The output from the audio encoder is denoted as $A_{E}\in\mathbb{R}^{d\times n}$, where d and n denote the embedding dimension and length of the audio (i.e. number of frames), respectively.
@@ G Q LPE LPH G Q LPE LPH
CTC [lugosch2018donut] 31.65 18.23 14.67 35.22 66.36 89.69 92.29 69.58
Attention [huang2021query] 14.75 49.13 28.74 41.95 92.09 50.13 78.74 62.65
Triplet [sacchi2019open] 35.6 38.72 32.75 44.36 71.48 66.44 63.53 54.88
CMCD [shin22_interspeech] 27.25 12.15 8.42 32.9 81.06 94.51 96.7 73.58
@@ II-C Pattern Extractor
It is based on the cross-attention mechanism [vaswani2017attention] which captures the temporal correlations between the audio and text embeddings. In this setup, the audio embedding $A_{E}$ functions as both the key and value, while the text embedding $T_{E}$ acts as the query. The output of the pattern extractor is the context vector which contains information about audio and text agreement.
@@ III-A Database
For training and evaluation, we used the Libriphrase [shin22_interspeech] dataset, which comprises short phrases with varying word lengths (1 to 4). The dataset was generated from the Librispeech corpus [panayotov2015librispeech]. The training set of Libriphrase is generated using train-clean-100 and train-clean-360 subsets, and the evaluation set is using train-others-500 subset. The evaluation set consists of 4391, 2605, 467, and 56 episodes of each word length respectively. Each episode has three positive and three negative pairs. The negative samples are further categorized into easy and hard based on Levenshtein distance [levenshtein1966binary], leading to the creation of the Libriphrase Easy (LPE) and Libriphrase Hard (LPH) datasets. Each sample in the dataset is represented by using three entities: audio, text, and a binary target value indicating 1 for a positive pair and 0 for a negative pair.
To comprehensively evaluate model performance, we extended our assessment beyond the Libriphrase dataset by incorporating the Google Speech Commands V1 dataset (G) [warden2018speech] and the Qualcomm Keyword Speech dataset (Q) [kim2019query]. The Google Speech Commands V1 dataset (G) encompasses speech recordings from 1,881 speakers, focusing on 30 small keywords. On the other hand, the Qualcomm Keyword Speech dataset (Q) comprises 4,270 utterances of four keywords spoken by 50 speakers.
@@ III-B Tacotron 2 Embeddings
In this study, the utilization of intermediate representation from Tacotron 2 is proposed to enhance performance, particularly in scenarios involving speech-keyword pairs with similar pronunciations. The intermediate representations are obtained using NVIDIA Tacotron 2 model [NVIDIATacotron2] which is pre-trained on the LJSpeech dataset [ito2017ljspeech]. The details of different intermediate representations that are used in this study are outlined in Table II. Additional insights into the Tacotron 2 architecture can be found in [shen2018natural]
@@ III-C Implementation Details
The training pipeline is structured as a binary classification task, wherein the model is tasked with classifying the similarity of input pairs {text, audio}. In audio and text encoders, weights are initialized through Xavier initialization, and Leaky ReLU is utilized as the activation function. A dropout of 0.2 is added after each layer in the encoders to prevent overfitting. The training process uses binary cross-entropy loss as the training criterion and employs the Adam optimizer [kingma2014adam] with default parameters for optimization. A fixed learning rate of $10^{-4}$ and a batch size of 128 is employed in the training process. The best-performing model was chosen based on its performance on the validation set. For training, we used four NVIDIA GeForce RTX 2080 Ti GPUs.